A gas at the temperature 250K is contained in a closed vessel. If the gas is heated through 1K, then the percentage increase in its pressure wil be...?

Grade:12th pass

1 Answers

Khimraj
5 years ago
PV = nRT
as V = consatnt
so 
P1/P2 = T1/T2
P1/P2 = 250/1250 = 1/5
P2 = 5P1
% change = ((5-1)/1)*100 = 400%
Hope it clears......................